OpenStudy (anonymous):

81 is 45% of what number?

OpenStudy (doc.brown):

\[\frac{81}{x}=\frac{45}{100}\]

OpenStudy (anonymous):

What Doc said

OpenStudy (doc.brown):

81:x::45:100

OpenStudy (anonymous):

I mean in general if you think about it, if you take something like let's say 50% of a number (let's use 60).....what you would generally do is multiply 60 by the decimal version of 50% which is 0.5. Using this logic, you can also say that if all that was given was "30 is 50% of what number" then you could take 30 and divide it by the decimal version of 50% and it should give you the answer of 60

OpenStudy (doc.brown):

\[x\times\frac{45}{100}=81\]

OpenStudy (doc.brown):

Well said @juanjohnguy

OpenStudy (anonymous):

Glad I can help.